Let x(y - 3) = 5 where x, y ∈ Integers then value of x + y is equal to-
(A) 6            (B) 9
(C) -6          (C) -3

Grade:11

2 Answers

Arun
6 years ago
Dear Student
 
If x and y are integers then y-3 will.also be a integer
Hence factor of 5 are = 1 *5 or 5 *1
Now
If x = 1, y-3 = 5
X = 1, y = 8
Hence x +y = 9
 
But if x = 5, y-3 = 1
X = 5, y = 4
Now x + y = 9
 
Hence ij every case answer is option B
 
Regards
Arun (askIITians forum expert)
downey zer
6 years ago
product of two integers = 5  
the only way this can happen is if the integers are either 5&1 or -5&-1
let x=1 and y-3=5   add these equations x+y=9
let x=5 and y-3=1 then x+y=9 again 
now let x=-1 and y-3=-5 then x+y=-3
similarly x=-5 and y-3=-1 will give x+y=-3
So answer is B,C